find the domain of the composite function f*g
f(x)=2/x+10,g(x)=60/x

Answer by edjones(8007)

f(x)=2/x+10,g(x)=60/x
f(g(x))=f(60/x)=2/((60/x)+10)
=2/((10x+60)/x)
=2x/10x+60
=x/5x+30
=x/5(x+6)
The domain of f*g is all real numbers except x=-6
.
Ed